OpenStudy (anonymous):

Someone Help plz!! Express the equation, θ = 5π/6 , in rectangular coordinates. (Use variables x and y.)

OpenStudy (anonymous):

Do you have your unit circle? You should be able to use that to find the answer

OpenStudy (anonymous):

@wrstlr3232 is it x^2+y^2=r^2?

OpenStudy (anonymous):

\[\sqrt{x ^{2}+y ^{2}}=r \]That's to find polar coordinates. You wouldn't need that for rectangular coordinates.

OpenStudy (anonymous):

You could use \[\frac{ 5\pi }{ 6 }\times \frac{ 180degrees }{ \pi }=150degrees\] That might make it easier to see where the point is on the unit circle.

OpenStudy (anonymous):

of course the coordinates in x and y\[x=r \cos \theta\]\[y=y \sin \theta\]and your angle \(\theta\) is in 2nd quadrant where x-ordinate must be negative (-) in sign, and y-ordinate is positive (+).